Project

General

Profile

Actions

20151201 » History » Revision 10

« Previous | Revision 10/33 (diff) | Next »
Boyd, Steven, 02 December 2015 11:35


Commissioning and Ops Scheduling meeting : 1/12/2015

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Focus coil progress and soak test
    • Decay Solenoid PSU Repair
    • Hydrogen/Vacuum system work
  • Upcoming activities : chaired by Victoria/Paul
    • This week's MICE Hall work plan
    • Focus coil operation
    • Hydrogen/Vacuum system installation and commissioning
    • Hydrogen alarm system work (requires roof access Tues--Fri)
    • Hall investigation for visual survey of X-ray leakage and reflection paths : Wed 10:30 am - 11:30 am (requires roof access)
    • Nitrogen system work (Tues--Thurs)
    • North-side handrail installation (Wed am)
  • Operations plan
    • Thursday 3rd (2 pm - 10 pm) : Noon run plan meeting / Close Hall / Beam tune / Prep data-taking / Take TOF 0 calibration (Q789 off) / FC off data / Open Hall
    • Friday 4th ( 12pm - 8 pm ) : 9 am Run plan meeting / Close Hall / Magnetic sweep / Ramp FC to 50 A Flip mode / FC on data-taking / Go to standby on the target overnight
    • Saturday 5th ( 9 am - 4 pm ) : 9 am Run Plan meeting / 10 am Ramp FC to 100 A / FC-on data taking / Go to standby on the target overnight
    • Sunday 6th ( 9 am - 4 pm ) : 9 am Run Plan meeting / 10 am Ramp to 150 A / FC-on data-taking / Go to standby on the target overnight
    • Monday 7th ( 9 am - end ) : 9 am run plan meeting / 10 am Ramp FC to 0 A / Change polarity / Ramp to -100 A / Take data / End and Open Hall
  • Questions
    • When can we start operations on Friday?
    • FC expert availability on Sat/Sun (drives when ramps can happen)
    • Hall activities on morning of Monday 7th
  • AOB
    • To-do list for Maintenance shutdown (January 2016)

Minutes

John C, Ryan, Melissa, Yordan, Celeste, Adam, Paolo, Ed, Durga, Victoria, Vicky, John, Mark, Steve

Review:

  • Focus coil ramped to 200 A over weekend and then to 225 A and then down. Looked stable.
  • Decay Solenoid PSU on transit. In Germany but has not arrive at FUG yet.
  • Hydrogen/Vacuum : Phil is in the Hall

Upcoming activities:

  • Focus coil operation went well so don't need further commissioning.
    John : did we lose any He on ramp-down.
    Mark : about 4-5 . We're at 68 now.
    Vicky : only boiled off during ramp-down
    John : should be OK over running period. 60% is the "safe" level.
  • Hydrogen/Vacuum system work on-going. Will take him until Thursday, leaving @ 12 on Thursday
  • John wants to work on the ceiling inside the Hall. Hope this will finish on Thursday.
  • Tim Stanley doing a walk-around on Wednesday mornig
  • Nitrogen system work : installation nitrogen
  • Handrail installed on Wednesday morning.
  • Two leaks has been found in the Hydrogen system. An external leak which can be fixed "easily". The major leak
    is on the inside of the bore in the magnet. Suspicion is that the leak is related to Indium o-rings but this can't
    be fixed from outside. We either need to change the medium or take the magnet apart to get at the leak. Plan to
    fix during January.
    Chris: can we do different absorbers now?
    Josef : yes. Could run methanol or water now if necessary. If we put in pressure we might collapse the end bore
    a bit.
    Mike : we could put methanol in there. It would have to be run past the safety committee.
    John : how would we get the methanol out.
    Mike : let it evaporate once we take the FC out.
    Ed : is there any possible risk in running in a mode we have not studied?
    Mike : need to think this through in the safety committee.
    Procedure : Steve W should drive this in consultation with the safety committee.
    Mike will get the ball rolling on getting this through the SC with the hope that we can run from the 14th to 18th.
    Chris will decide on methanol vs ethanol
    John will buy it in.
    Timescale to know if we're going to do this is one week. Next meeting is a decision point.
    John suggests removing the FC and putting a chunk of plastic in the beam?
    Ken doesn't think we should do methanol. Or water. Is there a gas which wouldn't freeze but would liquify? Argon?
    John says that you want a material which will have 5-10% of X0 in absorber? Pour in LN2 without the cooler?
    • Chris will look for "best" material and return a short list of possibilities by Thursday night.
    • Absorber team to look at late this week & into next week
  • Mike wants us to document the polarity change properly

AOB:

  • Josef: Do we want to press for the ECE coil at the magnet review?
    Victoria : would prefer ECE running

Updated by Boyd, Steven over 8 years ago · 10 revisions